# Universal Definitions of the Roman Factorial: Introduction to Foundational Functions and the Generalization Process

### v2

@@ 6.1 Introduction
$n!=\prod_{k=1}^{n}k\thinspace=1\cdot 2\cdot 3\cdots(n-2)(n-1)n\thinspace,\quad n\in\mathbb{N}\thinspace.$ (Eq. 6.1.1)
$n!=\prod_{k=0}^{n-1}(n-k)\thinspace=n(n-1)\cdots 3\cdot 2\cdot 1\thinspace,\quad n\in\mathbb{N}\thinspace.$ (Eq. 6.1.2)
@@
These procedures will be repeated for the falling product. The outcomes will be summarized, therefore concluding Part 1 of this study.
@@ 6.2 Rising product expression
Table Tbl. 6.2.1: Negative Roman factorials
@@
$\lfloor-3\rceil!=\cfrac{1}{2}=\cfrac{1}{-1}\cdot\cfrac{1}{-2}$
$\lfloor-4\rceil!=-\cfrac{1}{6}=\cfrac{1}{-1}\cdot\cfrac{1}{-2}\cdot\cfrac{1}{-3}$
$\lfloor-5\rceil!=\cfrac{1}{24}=\cfrac{1}{-1}\cdot\cfrac{1}{-2}\cdot\cfrac{1}{-3}\cdot\cfrac{1}{-4}$
@@
$\lfloor-3\rceil!=\cfrac{1}{-1}\cdot\cfrac{1}{-2}\cdot\cfrac{1}{-3}=-\cfrac{1}{6}$
$\lfloor-4\rceil!=\cfrac{1}{-1}\cdot\cfrac{1}{-2}\cdot\cfrac{1}{-3}\cdot\cfrac{1}{-4}=\cfrac{1}{24}$
$\lfloor-5\rceil!=\cfrac{1}{-1}\cdot\cfrac{1}{-2}\cdot\cfrac{1}{-3}\cdot\cfrac{1}{-4}\cdot\cfrac{1}{-5}=-\cfrac{1}{120}$
These values don’t match those in Tbl. 6.2.1. As previously noted, there’s an extra multiplication term at the end, which, if canceled, makes Eq. 6.2.1 accurate.
There are two approaches to address this issue: one is to adjust the upper limit of the product, subtracting 1 from it to multiply by one less term. The other involves multiplying the existing product by $n$, which cancels out the term causing the offset. Both methods work equally well, although the first idea utilizes the empty product [^14] For more information about the empty product, see Addendum 9.2. when $n=-1$.
We will proceed with the first approach for now, since it is more intuitive. Therefore, multiplying each negative Roman factorial by $n$ results in the following outcomes:
$\lfloor-3\rceil!=(-3)\cdot\cfrac{1}{-1}\cdot\cfrac{1}{-2}\cdot\cfrac{1}{-3}=\cfrac{1}{2}$
$\lfloor-4\rceil!=(-4)\cdot\cfrac{1}{-1}\cdot\cfrac{1}{-2}\cdot\cfrac{1}{-3}\cdot\cfrac{1}{-4}=-\cfrac{1}{6}$
$\lfloor-5\rceil!=(-5)\cdot\cfrac{1}{-1}\cdot\cfrac{1}{-2}\cdot\cfrac{1}{-3}\cdot\cfrac{1}{-4}\cdot\cfrac{1}{-5}=\cfrac{1}{24}$
@@
$\lfloor n\rceil!=n\cdot\prod_{k=1}^{-n}\cfrac{1}{-k}\enspace,\quad n\in\mathbb{Z}^{-}.$ (Eq. 6.2.2)
Notice the similarities between this expression and Eq. 6.1.1. We will shortly consider the case $n=0$ and form the piece-wise definition, although let’s firstly look at the analogous expression for the falling product.
@@ 6.3 Falling product expression
$\lfloor-3\rceil!=\cfrac{1}{2}=\cfrac{1}{-2}\cdot\cfrac{1}{-1}$
$\lfloor-4\rceil!=-\cfrac{1}{6}=\cfrac{1}{-3}\cdot\cfrac{1}{-2}\cdot\cfrac{1}{-1}$
$\lfloor-5\rceil!=\cfrac{1}{24}=\cfrac{1}{-4}\cdot\cfrac{1}{-3}\cdot\cfrac{1}{-2}\cdot\cfrac{1}{-1}$
@@
$\prod_{k=1}^{-n-1}\cfrac{1}{n+k}=\cfrac{1}{n+1}\cdot\cfrac{1}{n+2}\cdots\cfrac{1}{-2}\cdot\cfrac{1}{-1}\enspace,\quad n\in\mathbb{Z}^{-}.$ (Eq. 6.3.1)
While accurate and applicable to all negative integers, for $n=-1$, the use of the empty product is necessary again as the upper limit becomes smaller than the lower one.
To circumvent the empty product, we can multiply by one extra term and subsequently cancel it out, as demonstrated below:
$n\cdot\prod_{k=0}^{-n-1}\cfrac{1}{n+k}=\cfrac{n}{n}\cdot\cfrac{1}{n+1}\cdots\cfrac{1}{-2}\cdot\cfrac{1}{-1}\enspace,\quad n\in\mathbb{Z}^{-}.$ (Eq. 6.3.2)
@@
$\lfloor n\rceil!=n\cdot\prod_{k=0}^{-n-1}\cfrac{1}{n+k}\enspace,\quad n\in\mathbb{Z}^{-}.$ (Eq. 6.3.3)
@@
6.4 Generalization step 1: $|n|$
@@
$\lfloor n\rceil!=\begin{cases}\thinspace\displaystyle\prod_{k=1}^{n}k&,\enspace n\in\mathbb{N}\\[16.18028pt] n\cdot\displaystyle\prod_{k=1}^{-n}\cfrac{1}{-k}&,\enspace n\in\mathbb{Z}^{-}.\end{cases}$ (Eq. 6.4.1)
@@
$\lfloor n\rceil!=\begin{cases}\thinspace\displaystyle\prod_{k=0}^{n-1}(n-k)&,\enspace n\in\mathbb{N}\\[16.18028pt] n\cdot\displaystyle\prod_{k=0}^{-n-1}\cfrac{1}{n+k}&,\enspace n\in\mathbb{Z}^{-}.\end{cases}$ (Eq. 6.4.2)
@@
$n\in\mathbb{N}$ $\thinspace\displaystyle\prod_{k=1}^{n}k$ $\enspace\displaystyle\prod_{k=0}^{n-1}(n-k)$
@@
Since our definitions correctly describe the case of $n=0$, we can simply extend the domain of the positive-integer cases from $n\in\mathbb{N}$ to $n\in\mathbb{N}_{0}$. Therefore, the previous table is updated below:
@@
$n\in\mathbb{N}_{0}$ $\thinspace\displaystyle\prod_{k=1}^{n}k$ $\enspace\displaystyle\prod_{k=0}^{n-1}(n-k)$
@@
Next, we notice the upper limits of the products. For $n\in\mathbb{N}_{0}$, they are either $n$ or $n-1$, and for $n\in\mathbb{Z}^{-}$ they are $-n$ or $-n-1$.
Essentially, when $n$ becomes negative, a negative sign is applied to ensure that the upper limits remain positive numbers. As demonstrated in Subsection 4.6, we can replace occurrences of $n$ and $-n$ with $|n|$. Consequently, the table is modified as follows:
@@
$n\in\mathbb{N}_{0}$ $\thinspace\displaystyle\prod_{k=1}^{|n|}k$ $\enspace\displaystyle\prod_{k=0}^{|n|-1}(n-k)$
@@
Unlike Section 4, this section generalizes two definitions simultaneously. While it’s possible to separate the procedure for the rising product from the falling one, the steps are identical: combining the two Roman factorial product definitions into a single table allows us to present the generalization process concurrently.
6.5 Generalization step 2: $\Phi(n)$
@@
$\Phi(n)=\big{(}n+\Theta(n)\big{)}^{\xi^{\prime}(n)}=\begin{cases}1\thinspace,&n\in\mathbb{N}_{0}\\ n\thinspace,&n\in\mathbb{Z}^{-}.\end{cases}$ (Eq. 6.5.1)
Note that when $n$ is negative in Tbl. 6.4.3, the products are multiplied by a factor of $n$, which is absent in the products describing the $n\in\mathbb{N}_{0}$ case. This resembles the output pattern of $\Phi(n)$, which is $[n,1,1]$. Therefore, we can update Tbl. 6.4.3 as such:
@@
$n\in\mathbb{N}_{0}$ $\Phi(n)\cdot\displaystyle\prod_{k=1}^{|n|}k$ $\Phi(n)\cdot\displaystyle\prod_{k=0}^{|n|-1}(n-k)$
@@
6.6 Generalization step 3: $\theta(n)$
@@
It is evident that the sign of $k$, as well as the exponent, is represented by the same function. Additionally, it has the output pattern $[-1,1,1]$. A function that matches this requirement is $\theta(n)$:
$\theta(n)=\cfrac{\delta(n)}{|\delta(n)|}=\begin{cases}+1\thinspace,\quad n\in\mathbb{N}_{0}\\ -1\thinspace,\quad n\in\mathbb{Z}^{-}.\end{cases}$ (Eq. 6.6.1)
@@
$\theta(n)\cdot k^{\theta(n)}\rightarrow\big{(}k\cdot\theta(n)\big{)}^{\theta(n)}$
$\big{(}n-\theta(n)\cdot k\big{)}^{\theta(n)}\rightarrow\big{(}n-k\cdot\theta(n)\big{)}^{\theta(n)}$
@@
Rising product $\Phi(n)\cdot\displaystyle\prod_{k=1}^{|n|}\big{(}k\cdot\theta(n)\big{)}^{\theta(n)}$
Falling product $\Phi(n)\cdot\displaystyle\prod_{k=0}^{|n|-1}\big{(}n-k\cdot\theta(n)\big{)}^{\theta(n)}$
@@
Rising product $\prod_{k=1}^{|n|-\xi^{\prime}(n)}\big{(}k\cdot\theta(n)\big{)}^{\theta(n)}$
Falling product $\prod_{k=0}^{|n|-1-\xi^{\prime}(n)}\big{(}n-k\cdot\theta(n)\big{)}^{\theta(n)}$
@@ 7.1 Summary
  • Section 2: We analyzed the factorial and presented the Roman factorial, an expansion for negative integers.
@@
  • Section 4: We employed the F.F. shown in the previous section to rewrite the Roman factorial definition concisely.
@@ 7.2 Results
$\lfloor n\rceil!=\begin{cases}n!&,\enspace n\in\mathbb{N}_{0}\\ \cfrac{(-1)^{-n-1}}{(-n-1)!}&,\enspace n\in\mathbb{Z}^{-},\end{cases}$ (Eq. 7.2.1)
@@
$n!=n(n-1)!\thinspace,\quad 0!=1\thinspace,\quad n\in\mathbb{N}\thinspace.$ (Eq. 7.2.2)
@@
$\lfloor n\rceil!=\eta(n)\cdot{\big{[}|n|-\xi^{\prime}(n)\big{]}!}^{\theta(n)},\quad n\in\mathbb{Z}\thinspace,$ (Eq. 7.2.3)
@@
$n!=n(n-1)!\thinspace,\quad 0!=1\thinspace,\quad n\in\mathbb{N}\thinspace.$ (Eq. 7.2.4)
@@
$\lfloor n\rceil!=\begin{cases}n\lfloor n-1\rceil!&,\enspace n\in\mathbb{N}\\[3.68027pt] \cfrac{\lfloor n+1\rceil!}{n+1}&,\enspace n\in\mathbb{Z}^{-}\setminus\{-1\}\thinspace,\end{cases}$ (Eq. 7.2.5)
@@
$\lfloor n\rceil!=\big{(}n+\xi^{\prime}(n)\big{)}^{\theta(n)}\lfloor n-\theta(n)\rceil!\thinspace,\enspace n\in\mathbb{Z}\setminus\{0,\enspace-1\}\thinspace,$ (Eq. 7.2.7)
@@
$n\in\mathbb{N}_{0}$ $\thinspace\displaystyle\prod_{k=1}^{n}k$ $\enspace\displaystyle\prod_{k=0}^{n-1}(n-k)$
@@
Table Tbl. 7.2.1: Roman factorial as a rising or falling product
@@
Rising product $\Phi(n)\cdot\displaystyle\prod_{k=1}^{|n|}\big{(}k\cdot\theta(n)\big{)}^{\theta(n)}$
Falling product $\Phi(n)\cdot\displaystyle\prod_{k=0}^{|n|-1}\big{(}n-k\cdot\theta(n)\big{)}^{\theta(n)}$
Table Tbl. 7.2.2: Roman factorial as a rising or falling product (generalized)
In the next part, we will examine the double factorial and follow a similar procedure. We will investigate an extension into negative integers, we will find recursive as well as non-recursive piece-wise definitions of the expanded double factorial and we will unify their cases. This process will be repeated for the triple factorial and in general for all factorials of higher order.

### v7

@@ 6.1 Introduction
This section will focus on generalizing the $\prod$-product definitions of the Roman factorial, in the same manner as it was done in Section 4.
@@
$n!=\prod_{k=1}^{n}k\,=1\cdot 2\cdot 3\cdots(n-2)(n-1)n\,,\quad n\in\mathbb{Z}^{+},$ (Eq. 2.1.1)
$n!=\prod_{k=0}^{n-1}(n-k)\,=n(n-1)\cdots 3\cdot 2\cdot 1\,,\quad n\in\mathbb{Z}^{+}.$ (Eq. 2.1.2)
@@
These procedures will be repeated for the falling product. The outcomes will then be summarized, therefore concluding Part 1 of this study.
@@ 6.2 Rising product expression
Table Tbl. 2.2.1: Negative Roman factorials
@@
$\lfloor-3\rceil!=\cfrac{1}{2}=\cfrac{1}{-1}\cdot\cfrac{1}{-2}\;,$
$\lfloor-4\rceil!=-\cfrac{1}{6}=\cfrac{1}{-1}\cdot\cfrac{1}{-2}\cdot\cfrac{1}{-3}\;,$
$\lfloor-5\rceil!=\cfrac{1}{24}=\cfrac{1}{-1}\cdot\cfrac{1}{-2}\cdot\cfrac{1}{-3}\cdot\cfrac{1}{-4}\;.$
@@
$\lfloor-3\rceil!=\cfrac{1}{-1}\cdot\cfrac{1}{-2}\cdot\cfrac{1}{-3}=-\cfrac{1}{6}\>,$
$\lfloor-4\rceil!=\cfrac{1}{-1}\cdot\cfrac{1}{-2}\cdot\cfrac{1}{-3}\cdot\cfrac{1}{-4}=\cfrac{1}{24}\>,$
$\lfloor-5\rceil!=\cfrac{1}{-1}\cdot\cfrac{1}{-2}\cdot\cfrac{1}{-3}\cdot\cfrac{1}{-4}\cdot\cfrac{1}{-5}=-\cfrac{1}{120}\>.$
These values don’t match those in Tbl. 2.2.1. As previously noted, there’s an extra multiplication term at the end, which, if canceled, makes Eq. 6.2.1 accurate.
There are two approaches to address this issue: one is to adjust the upper limit of the product, subtracting 1 from it to multiply by one less term. The other involves multiplying the existing product by $n$, which cancels out the term causing the offset. Both methods work equally well, although the first idea utilizes the empty product [^17] For more information about the empty product, check Addendum 9.2. when $n=-1$.
We will proceed with the second approach for now. Therefore, multiplying each negative Roman factorial by $n$ results in the following outcomes:
$\lfloor-3\rceil!=(-3)\cdot\cfrac{1}{-1}\cdot\cfrac{1}{-2}\cdot\cfrac{1}{-3}=\cfrac{1}{2}\>,$
$\lfloor-4\rceil!=(-4)\cdot\cfrac{1}{-1}\cdot\cfrac{1}{-2}\cdot\cfrac{1}{-3}\cdot\cfrac{1}{-4}=-\cfrac{1}{6}\>,$
$\lfloor-5\rceil!=(-5)\cdot\cfrac{1}{-1}\cdot\cfrac{1}{-2}\cdot\cfrac{1}{-3}\cdot\cfrac{1}{-4}\cdot\cfrac{1}{-5}=\cfrac{1}{24}\>.$
@@
$\lfloor n\rceil!=n\cdot\prod_{k=1}^{-n}\cfrac{1}{-k}\;,\quad n\in\mathbb{Z}^{-}.$ (Eq. 6.2.2)
Notice the similarities between this expression and Eq. 2.1.1. We will eventually consider the case of $n=0$ and form the piece-wise definition, although let’s firstly look at the analogous expression for the falling product.
@@ 6.3 Falling product expression
$\lfloor-3\rceil!=\cfrac{1}{2}=\cfrac{1}{-2}\cdot\cfrac{1}{-1}\>,$
$\lfloor-4\rceil!=-\cfrac{1}{6}=\cfrac{1}{-3}\cdot\cfrac{1}{-2}\cdot\cfrac{1}{-1}\>,$
$\lfloor-5\rceil!=\cfrac{1}{24}=\cfrac{1}{-4}\cdot\cfrac{1}{-3}\cdot\cfrac{1}{-2}\cdot\cfrac{1}{-1}\>.$
@@
$\prod_{k=1}^{-n-1}\cfrac{1}{n+k}=\cfrac{1}{n+1}\cdot\cfrac{1}{n+2}\cdots\cfrac{1}{-2}\cdot\cfrac{1}{-1}\;,\quad n\in\mathbb{Z}^{-}.$ (Eq. 6.3.1)
While accurate and applicable to all negative integers, for $n=-1$, the use of the empty product is necessary again as the upper limit becomes smaller than the lower one. It is preferable to have a definition with the same lower product limit $k=0$ as in Eq. 2.1.2. So, we can multiply by one extra term and subsequently cancel it out, as demonstrated below:
$n\cdot\prod_{k=0}^{-n-1}\cfrac{1}{n+k}=\cfrac{n}{n}\cdot\cfrac{1}{n+1}\cdots\cfrac{1}{-2}\cdot\cfrac{1}{-1}\;,\quad n\in\mathbb{Z}^{-}.$ (Eq. 6.3.2)
@@
$\lfloor n\rceil!=n\cdot\prod_{k=0}^{-n-1}\cfrac{1}{n+k}\;,\quad n\in\mathbb{Z}^{-}.$ (Eq. 6.3.3)
@@
6.4 $\prod$-product definition step 1: $|n|$
@@
$\lfloor n\rceil!=\begin{cases}\,\displaystyle\prod_{k=1}^{n}k&,\enspace n\in\mathbb{Z}^{+}\\[16.18028pt] n\cdot\displaystyle\prod_{k=1}^{-n}\cfrac{1}{-k}&,\enspace n\in\mathbb{Z}^{-}.\end{cases}$ (Eq. 6.4.1)
@@
$\lfloor n\rceil!=\begin{cases}\,\displaystyle\prod_{k=0}^{n-1}(n-k)&,\enspace n\in\mathbb{Z}^{+}\\[16.18028pt] n\cdot\displaystyle\prod_{k=0}^{-n-1}\cfrac{1}{n+k}&,\enspace n\in\mathbb{Z}^{-}.\end{cases}$ (Eq. 6.4.2)
@@
$n\in\mathbb{Z}^{+}$ $\,\displaystyle\prod_{k=1}^{n}k$ $\enspace\displaystyle\prod_{k=0}^{n-1}(n-k)$
@@
Since our definitions correctly describe the case of $n=0$, we can simply extend the domain of the $n\in\mathbb{Z}^{+}$ cases to $n\in\mathbb{Z}^{+}_{0}$. Therefore, the previous table is updated below:
@@
$n\in\mathbb{Z}^{+}_{0}$ $\,\displaystyle\prod_{k=1}^{n}k$ $\enspace\displaystyle\prod_{k=0}^{n-1}(n-k)$
@@
Next, we notice the upper limits of the products. For $n\in\mathbb{Z}^{+}_{0}$, they are either $n$ or $n-1$, and for $n\in\mathbb{Z}^{-}$ they are $-n$ or $-n-1$.
Essentially, when $n$ becomes negative, a negative sign is applied to ensure that the upper limits remain positive numbers. As seen in Subsection 4.6, we can replace occurrences of $n$ and $-n$ with $|n|$. Consequently, the table is modified as follows:
@@
$n\in\mathbb{Z}^{+}_{0}$ $\,\displaystyle\prod_{k=1}^{|n|}k$ $\enspace\displaystyle\prod_{k=0}^{|n|-1}(n-k)$
@@
Unlike Section 4, this section generalizes two definitions simultaneously. While it’s possible to separate the procedure for the rising product from the falling one, the steps are identical: combining the two Roman factorial product definitions into a single table allows us to present the generalization process concurrently, unifying the expressions top to bottom.
6.5 $\prod$-product definition step 2: $\Phi(n)$
@@
$\Phi(n)=\big{(}n+\Theta(n)\big{)}^{\xi^{\prime}(n)}=\begin{cases}1\,,&n\in\mathbb{Z}^{+}_{0}\\ n\,,&n\in\mathbb{Z}^{-}.\end{cases}$ (Eq. 5.6.3)
Note that when $n$ is negative in Tbl. 6.4.3, the products are multiplied by a factor of $n$, which is absent in the products describing the $n\in\mathbb{Z}^{+}_{0}$ case. This resembles the output pattern of $\Phi(n)$, which is $[n,1,1]$. Therefore, we can update Tbl. 6.4.3 as such:
@@
$n\in\mathbb{Z}^{+}_{0}$ $\Phi(n)\cdot\displaystyle\prod_{k=1}^{|n|}k$ $\Phi(n)\cdot\displaystyle\prod_{k=0}^{|n|-1}(n-k)$
@@
6.6 $\prod$-product definition step 3: $\theta(n)$
@@
It is evident that the sign of $k$, as well as the exponent, is represented by the same function. That function has the output pattern $[-1,1,1]$ and a F.F. that matches this requirement is $\theta(n)$:
$\theta(n)=\cfrac{\delta(n)}{|\delta(n)|}=\begin{cases}+1\,,\quad n\in\mathbb{Z}^{+}_{0}\\ -1\,,\quad n\in\mathbb{Z}^{-}.\end{cases}$ (Eq. 3.3.2)
@@
$\theta(n)\cdot k^{\theta(n)}\rightarrow\big{(}k\,\theta(n)\big{)}^{\theta(n)}$
$\big{(}n-\theta(n)\cdot k\big{)}^{\theta(n)}\rightarrow\big{(}n-k\,\theta(n)\big{)}^{\theta(n)}$
@@
Rising product $\Phi(n)\cdot\displaystyle\prod_{k=1}^{|n|}\big{(}k\,\theta(n)\big{)}^{\theta(n)}$
Falling product $\Phi(n)\cdot\displaystyle\prod_{k=0}^{|n|-1}\big{(}n-k\,\theta(n)\big{)}^{\theta(n)}$
@@
Rising product $\prod_{k=1}^{|n|-\xi^{\prime}(n)}\big{(}k\,\theta(n)\big{)}^{\theta(n)}$
Falling product $\prod_{k=0}^{|n|-1-\xi^{\prime}(n)}\big{(}n-k\,\theta(n)\big{)}^{\theta(n)}$
@@
These alternative $\prod$-products would arise if we chose a different approach in Subsections 6.2 and 6.3, when we chose to multiply the existing products by $n$ instead of altering their upper limits.
Here we see that there is a F.F. in the limits of the $\prod$-products, a result we avoided for two reasons: simplicity of the expression and easier generalization in later parts of this study. The last reason was not explained here because of lacking context, but it will become apparent in the next part. There, the same process of generalization will apply to the multifactorial and its expansions into negative integers.
@@ 7.1 Summary
  • Section 2: We analyzed the factorial and presented the Roman factorial.
@@
  • Section 4: We employed the F.F. shown in the previous section to rewrite the Roman factorial definition concisely, in its original formulation as well as its recursive form.
@@
$\delta(n)$ $=\lfloor n\rfloor+0.5$ [,+,+[-,+,+][ - , + , + ] (Eq. 3.2.3)
$\theta(n)$ $=\cfrac{\delta(n)}{\left|\delta(n)\right|}$ $[-1,1,1]$ (Eq. 3.3.2)
$\xi(n)$ $=\cfrac{1+\theta(n)}{2}$ $[0,1,1]$ (Eq. 3.4.2)
$\xi^{\prime}(n)$ $=\cfrac{1-\theta(n)}{2}$ $[1,0,0]$ (Eq. 3.5.1)
@@
$\Theta(n)$ $=\xi(n)\cdot\xi(-n)$ $[0,1,0]$ (Eq. 5.2.2)
$Q(n)$ $=\theta(n)-\Theta(n)$ $[-1,0,1]$ (Eq. 5.3.5)
$Q^{\prime}(n)$ $=1-\Theta(n)$ $[1,0,1]$ (Eq. 5.4.2)
$\Psi(n)$ $=n+\Theta(n)$ $[n,1,n]$ (Eq. 5.5.2)
$\Phi(n)$ $={\Psi(n)}^{\xi^{\prime}(n)}$ $[n,1,1]$ (Eq. 5.6.3)
@@ 7.2 Results
$\lfloor n\rceil!=\begin{cases}n!&,\enspace n\in\mathbb{Z}^{+}_{0}\\ \cfrac{(-1)^{-n-1}}{(-n-1)!}&,\enspace n\in\mathbb{Z}^{-},\end{cases}$ (Eq. 2.2.1)
@@
$n!=n(n-1)!\,,\quad 0!=1\,,\quad n\in\mathbb{Z}^{+}.$ (Eq. 2.1.3)
@@
$\lfloor n\rceil!=\eta(n)\cdot\big{(}|n|-\xi^{\prime}(n)\big{)}!^{\,\theta(n)}\,,\quad n\in\mathbb{Z}\,,$ (Eq. 4.7.3)
@@
$n!=n(n-1)!\,,\quad 0!=1\,,\quad n\in\mathbb{Z}^{+}.$ (Eq. 2.1.3)
@@
$\lfloor n\rceil!=\begin{cases}n\lfloor n-1\rceil!&,\enspace n\in\mathbb{Z}^{+}\\[3.68027pt] \cfrac{\lfloor n+1\rceil!}{n+1}&,\enspace n\in\mathbb{Z}^{-}\setminus\{-1\}\,,\end{cases}$ (Eq. 2.2.3)
@@
$\lfloor n\rceil!=\big{(}n+\xi^{\prime}(n)\big{)}^{\theta(n)}\lfloor n-\theta(n)\rceil!\,,\quad n\in\mathbb{Z}\setminus\{0,\enspace-1\}\,,$ (Eq. 4.3.2)
@@
$n\in\mathbb{Z}^{+}_{0}$ $\,\displaystyle\prod_{k=1}^{n}k$ $\enspace\displaystyle\prod_{k=0}^{n-1}(n-k)$
@@
Table Tbl. 6.4.2: Roman factorial as a rising or falling product
@@
Rising product $\Phi(n)\cdot\displaystyle\prod_{k=1}^{|n|}\big{(}k\,\theta(n)\big{)}^{\theta(n)}$
Falling product $\Phi(n)\cdot\displaystyle\prod_{k=0}^{|n|-1}\big{(}n-k\,\theta(n)\big{)}^{\theta(n)}$
In the next part, we will examine the double factorial and follow a similar procedure. We will investigate an extension into negative integers, we will find recursive as well as non-recursive piece-wise definitions of the expanded double factorial and we will unify their cases. This process will be repeated for the triple factorial and in general for all factorials of higher orders.